OpenStudy (anonymous):

Graph the function below. Determine the domain, range, and horizontal asymptote. Please show all of your work. f(x) = 3 - (1/2)^x

OpenStudy (anonymous):

ok we go slow domain: all real numbers since you can raise 1/2 to any power you like

OpenStudy (anonymous):

range: no matter what x is \[(\frac{1}{2})^x>0\] and so \[3-(\frac{1}{2})^x<3\] and your range is \[(-\infty, 3)\]

OpenStudy (anonymous):

since that is your range, you know horizontal asymptote is \[y=3\] and here is a decent picture

OpenStudy (anonymous):

i dont see the picture. lol

OpenStudy (anonymous):

so what would the domain be?